# Parameter-Efficient Subspace Optimization for LLM Fine-Tuning

### v1

@@ 4.3 Multi-Epoch Low-Rank Analysis
To demonstrate the effectiveness of subspace exploration, we extend T5-base fine-tuning on SST-2 from one epoch (Section 4.1) to four. This longer schedule enables more thorough exploration and reduces the intrinsic low-rank bottleneck. As shown in Table 6, PESO-LoRA-R with $r=2$ consistently outperforms standard LoRA even at higher ranks ($r=4,8$), showing that it overcomes the low-rank limitation, achieves full-parameter optimality, and delivers stronger performance even under highly restricted memory budgets.
Table 6: Performance of fine-tuned T5-base (4 epochs) on the SST-2 dataset. Results are reported as accuracy (%) over 3 runs. Bold and underline indicate the highest and second-highest accuracies.
[…]
black
Method Epoch 1 Epoch 2 Epoch 3 Epoch 4
@@ 5 Convergence Analysis
In this section, we establish convergence of PESO with the full gradient restart. For general nonconvex losses, the optimality measure $\mathbb{E}\|G_{k}\|$ converges to zero up to controlled inexactness. We begin by stating the regularity assumptions.
@@
The loss $\ell$ is nonconvex, bounded from below, and has $L$-Lipschitz gradients.
@@
Stochastic gradients $\widetilde{G}_{k}$ of the full gradient $G_{k}$ satisfy $\mathbb{E}(\widetilde{G}_{k})=G_{k}$ and there exists $C>0$ such that $\mathbb{V}(\widetilde{G}_{k})\leq C$.
@@
There exists a sequence $\{\delta_{k}\}\geq 0$ such that $dist(G_{k},\mathcal{S}_{k})\leq\delta_{k}$ for $k$ where full gradient restart is implemmented. Furthermore, $\lim_{k\to\infty}\delta_{k}<\infty$.
Assumption 5.
Opt and UpdateSubspace generate the updates satisfying $\mathbb{E}[\ell(W_{k})]\leq\mathbb{E}[\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))]+C_{k}$ and $\mathbb{E}[\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))]\leq\mathbb{E}[\ell(W_{k-1})]+C_{k}$ where $\sum_{k}|C_{k}|<\infty$.
Assumptions 13 are standard in stochastic optimization; see, e.g., [bottou2018optimization]. Assumption 4 requires the subspace at full gradient restart to be sufficiently aligned with $G_{k}$, allowing approximation errors, e.g., from low-rank SVD. Assumption 5 is mild and holds, for example, when both Opt and UpdateSubspace use SGD with diminishing learning rates; see Appendix G.
We now state the main convergence result; the proof and its deterministic counterpart are deferred to Appendix G. Importantly, the result holds for any choice of UpdateSubspace, whether warm-start or restart, and at any frequency. In particular, Algorithm 1 may combine different UpdateSubspace strategies at varying frequencies, and the guarantee remains valid.
@@
Suppose all assumptions hold. With full gradient restart, the iterates $\{W_{k}\}$ generated by Algorithm 1 satisfy $\liminf_{k\to\infty}\mathbb{E}[\|G_{k}\|]\leq\lim_{k\to\infty}\delta_{k}$.
Remark 1.
If $\mathcal{S}_{k}^{\text{FG}}$ is chosen such that $G_{k}\in\mathcal{S}_{k}^{\text{FG}}$, then $\delta_{k}=0$ and the optimality measure converges to zero. In PESO-LoRA-R, $\mathcal{S}_{k}^{\text{FG}}$ is the top-$r$ SVD subspace of $\widetilde{G}_{k}$, so $\delta_{k}$ reflects both SVD truncation error and a noise term of order $O(\sqrt{C})$. When $G_{k}$ is effectively low rank and variance-reduction (e.g., online PCA or EMA) is used, $\lim_{k\to\infty}\delta_{k}$ can be made small.
Remark 2.
As a simple instantiation of PESO, we note that in the deterministic setting where $\mathcal{S}_{k}^{\text{FG}}$ is chosen as the top-$r$ SVD subspace of $G_{k}$, as in PESO-LoRA-R, the bias term in Theorem 5.1 disappears. In this case, we can establish exact asymptotic convergence, $\liminf_{k\to\infty}\mathbb{E}[\|G_{k}\|]=0$, i.e., PESO-LoRA-R converges to a first-order stationary point; see full details in Appendix G.3.
@@ Appendix A Synthetic Example of LoRA Deficiency
One critical limitation in the literature is the absence of convergence guarantees toward valid optimality conditions of (1). Most existing works establish convergence only with respect to the low-dimensional parameters—such as the factors $A$ and $B$ in LoRA—but do not address convergence with respect to the full-parameter matrix $W$. For instance, [jiang2024unified] shows that $\nabla_{A}\ell(W_{0}+A_{k}B_{k})$ and $\nabla_{B}\ell(W_{0}+A_{k}B_{k})$ vanish as $k\to\infty$, while leaving the behavior of $\nabla_{W}\ell$ uncharacterized. This gap is not merely technical: it highlights a fundamental deficiency of PEFT methods compared to standard full-parameter training. In fact, the optimal loss attained by LoRA can be arbitrarily worse than the true optimal loss of (1). To illustrate this, consider the following simple synthetic example in matrix optimization:
@@ Appendix B Review on Subspace Minimization
It is worth noting that Algorithm 1 is essentially equivalent to the classical two-loop subspace minimization scheme [conn1994iterated], summarized in Algorithm 4.
@@
1:Initialization $W_{0}\in\mathbb{R}^{m\times n}$, $\xi_{0}\in\mathbb{R}^{d}$, $\mathcal{M}_{0}$, an algorithmic subroutine UpdateSubspace, an optimizer Opt, frequency $K$.
@@
4:  Update the subspace $(\mathcal{M}_{k},\widetilde{W}_{k})\leftarrow\texttt{UpdateSubspace}(\mathcal{M}_{k-1},\widetilde{W}_{k-1})$.
@@ Appendix C Projected Subspace and Memory Efficiency
One-sided projected subspaces in PESO can offer stronger memory efficiency than LoRA. This idea is exemplified by GaLore [zhao2024galore], which we now place in the PESO framework. GaLore requires memory of order $mn+mr+2nr$ (assuming $m\leq n$), compared to LoRA’s $mn+3mr+3nr$.
@@
which matches the classical projected subspace descent step [kozak2019stochastic].
@@ Appendix D Implementation of PESO-LoRA-R
We discuss several implementation details of PESO-LoRA-R that are critical for practical stability and performance.
@@ D.1 Smoothing the Subspaces
To mitigate this, we adopt an Exponential Moving Average (EMA) of old and new subspaces, similar in spirit to how Adam [kingma2014adam] stabilizes noisy updates. However, this is nontrivial in PESO-LoRA-R (Algorithm 2), since the pre-restart adapters $(A_{k},B_{k})$—evolved through Adam dynamics—may differ significantly in scale and coordinates from the restarted pair $(U_{k}\sqrt{\Lambda_{k}},\sqrt{\Lambda_{k}}V_{k})$ obtained from rank-$r$ SVD. A naive EMA would mismatch these terms and discard valuable exploration information.
@@
with parameter $\tau_{2}$. Since this merges the scaling of $A$ and $B$, we refactorize $S_{\mathrm{new}}$ using polar decomposition [glentis2025scalable]:
@@ Appendix E Detailed Version of Algorithms
In this section, we provide detailed versions of PESO-LoRA-R (Algorithm 2) and PESO-LoRA-T (Algorithm 3).
Algorithm 5 PESO-LoRA-R: PESO with LoRA and Subspace ExploRation
1:Pre-trained parameters $W_{0}\in\mathbb{R}^{m\times n}$, frequency $K$, scale parameter $\gamma$, learning rate $\eta$, AdamW hyperparameters $(\beta_{1},\beta_{2},\varepsilon,\lambda)$.
@@
17:  $A_{k}\leftarrow A_{k-1}-\eta\lambda A_{k-1}-\eta\,\widehat{m}_{A}/(\sqrt{\widehat{v}_{A}}+\varepsilon)$.
18:  $B_{k}\leftarrow B_{k-1}-\eta\lambda B_{k-1}-\eta\,\widehat{m}_{B}/(\sqrt{\widehat{v}_{B}}+\varepsilon)$.
@@
Algorithm 6 PESO-LoRA-T: PESO with LoRA and Subspace ExploiTation
1:Pretrained weights $W_{0}\in\mathbb{R}^{m\times n}$, initial subspace matrices $U_{0}\in\mathbb{R}^{m\times r}$, $V_{0}\in\mathbb{R}^{r\times n}$, initial coordinate $\xi_{0}\in\mathbb{R}^{r}$, frequency $K$, learning rate $\eta$, AdamW hyperparameters $(\beta_{1},\beta_{2},\varepsilon,\lambda)$.
@@
13:   $U_{k}\leftarrow U_{k-1}-\eta\lambda U_{k-1}-\eta\,\widehat{m}_{U}/(\sqrt{\widehat{v}_{U}}+\varepsilon)$.
14:   $V_{k}\leftarrow V_{k-1}-\eta\lambda V_{k-1}-\eta\,\widehat{m}_{V}/(\sqrt{\widehat{v}_{V}}+\varepsilon)$.
@@
19:  $\xi_{k}\leftarrow\xi_{k-1}-\eta\lambda\xi_{k-1}-\eta\,\widehat{m}_{\xi}/(\sqrt{\widehat{v}_{\xi}}+\varepsilon)$.
@@ F.1 Natural Language Understanding: Hyperparameter Settings
In Section 4.1, we present the results of our methods and various LoRA-based algorithms on natural language understanding tasks, following the prompt tuning configuration of [wang2024loraga]. The general hyperparameter settings are kept consistent across all algorithms which are shown in Table 7. To ensure a fair comparison, we follow [zhang2025one] and tune the learning rates via grid search over $\{1\times 10^{-4},2\times 10^{-4},5\times 10^{-4},1\times 10^{-3}\}$. Additionally, following the choices of [zhang2025one], the scale parameters for LoRA-One are set to be {128, 16, 128, 128, 64} for {MNLI, SST-2, CoLA, QNLI, MRPC}.
@@ F.2.1 Comparison with LoRA-based methods
For natural language generation tasks in Section 4.2, we follow the configuration of prompt tuning and strategy of hyperparameter tuning in [zhang2025one] to ensure fair comparison. We search the best learning rate over $\{5\times 10^{-4},\,2\times 10^{-4},\,1\times 10^{-4},\,5\times 10^{-5},\,2\times 10^{-5},\,1\times 10^{-5}\}$, and the general hyperparameter setting is summarized in Table 8. Additionally, following the choice of [zhang2025one], the scale parameters are tuned within $\{16,32,64,128\}$ for LoRA-GA, LoRA-One and our methods to achieve the best performances.
@@ F.4 Multi-Epoch Low-Rank Analysis: Hyperparameter Settings
Appendix G Proofs
In this section, we provide the proofs of the theoretical results stated in Section 5. For completeness, we begin with the deterministic case, i.e., when the gradients $G_{k}$ accessed by PESO (Algorithm 1) are exact, without stochastic noise. We then prove Theorem 5.1, which considers the stochastic setting where only noisy gradients $\widetilde{G}_{k}$ are available, satisfying Assumption 2.
@@ G.1 Deterministic Case
We first state the deterministic counterpart of Assumption 5:
@@
Opt and UpdateSubspace generate the updates satisfying $\ell(W_{k})\leq\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))$ and $\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))\leq\ell(W_{k-1})$ for all $k=1,2,\cdots$.
Assumption 6 requires that both Opt and UpdateSubspace act as descent methods, ensuring that the loss is non-increasing. This is the deterministic analogue of Assumption 5, and it holds, for example, when both are implemented as gradient descent with step size $\alpha\leq 1/L$, where $L$ is the Lipschitz constant from Assumption 1. In particular, one can take UpdateSubspace to be a warm-start step that performs gradient descent on the subspace parameters with respect to the original loss function. This is formalized below in a standard result from the optimization literature see, e.g., [nesterov2018lectures].
@@
Let Opt and UpdateSubspace be gradient descent schemes on $\ell$ with constant learning rate $\alpha\leq 1/L$. Then Assumption 6 holds.
@@
Suppose Assumptions 1, 4, and 6 hold. With full gradient restart enabled and learning rate $\eta_{k}=\tfrac{1}{L}$, the iterates $\{W_{k}\}$ generated by Algorithm 1 satisfy $\liminf_{k\to\infty}\|G_{k}\|\leq\lim_{k\to\infty}\delta_{k}$.
@@
Performing a full gradient restart as UpdateSubspace, as discussed in (8), is thus equivalent to moving from $W_{k-1}$ to $\widehat{W}_{k}$. It follows that
@@
Importantly, (15) remains valid regardless of how frequently other types of UpdateSubspace are applied between full gradient restarts, since all updates preserve the descent property by Assumption 6. In particular, whenever $\mathcal{M}_{k}$ is updated without a full gradient restart, we have

### v2

@@ 4.3 Multi-Epoch Low-Rank Analysis
To highlight the effect of subspace exploration, we extend T5-base fine-tuning on SST-2 from one epoch (Section 4.1) to four. The longer schedule allows more thorough exploration and mitigates the intrinsic low-rank bottleneck. As shown in Table 4, PESO-LoRA-R with $r=2$ consistently outperforms standard LoRA even with larger ranks ($r=4,8$), indicating that it alleviates the low-rank limitation and achieves stronger performance under tightly restricted memory budgets.
@@ 5 Convergence Analysis
In this section we establish convergence guarantees for PESO-LoRA-R, a practical instantiation of PESO. A more general asymptotic convergence result for PESO is provided in Appendix G. We consider a nonconvex objective $\ell(W)$ and make the following standard assumptions.
@@
The loss $\ell$ is bounded below and has $L$-Lipschitz continuous gradients.
@@
Stochastic gradients $\widetilde{G}_{k}$ of the full gradient $G_{k}=\nabla\ell(W_{k})$ are accessible and satisfy $\mathbb{E}(\widetilde{G}_{k})=G_{k}$. There exists $C>0$ such that $\mathbb{V}(\widetilde{G}_{k})\leq C$.
We analyze PESO-LoRA-R (Algorithm 2), where the inner optimizer is set as Adam. Below we state a standard regularity assumption on the low-rank objective $f(A,B)=\ell(W+AB)$ optimized by Adam.
@@
At each iteration, PESO-LoRA-R applies Adam to the low-rank objective $f(A,B)=\ell(W+AB)$. We assume for $f$: (i) $\nabla f$ is Lipschitz continuous, (ii) stochastic gradients are unbiased, and (iii) stochastic gradient variance is bounded.
Let $P_{r}(G_{k})$ denote the rank-$r$ truncated SVD of $G_{k}$. Since $P_{r}(\cdot)$ is nonlinear, in general $\mathbb{E}[P_{r}(\widetilde{G}_{k})]\neq P_{r}(\mathbb{E}[\widetilde{G}_{k}])$. Our analysis only requires that this projection-induced bias is asymptotically negligible in a stepsize-weighted sense.
@@
At each restart step $(k-1)\bmod K=0$, the estimated subspace satisfies $\|\mathbb{E}[P_{r}(\widetilde{G}_{k})]-P_{r}(G_{k})\|_{F}\leq\varepsilon_{k},\sum_{k}\eta_{k}\varepsilon_{k}^{2}<\infty.$
Remark 1.
This assumption holds in several standard regimes in subspace optimization literature. It holds trivially in the full-batch setting ($\varepsilon_{k}=0$) and is satisfied in the large-batch regime via concentration bounds $\varepsilon_{k}=\mathcal{O}(|\mathcal{B}|^{-1/2})$; see, e.g., (He et al., 2024). Furthermore, using independent randomized subspaces (Chen et al., 2025) yields $\varepsilon_{k}=0$ due to conditional unbiasedness.
@@
Under Assumptions 14 and following the hyperparameter choices from Appendix H, the iterates generated by PESO-LoRA-R satisfy
$\min_{0\leq k\leq T-1}\mathbb{E}\|\nabla\ell(W_{k})\|_{2}^{2}=\mathcal{O}\!\left(1/{\sqrt{T}}\right).$
@@ 6 Conclusions and Limitations
Acknowledgments
We would like to thank Professor Jorge Nocedal for his invaluable insights and discussions, which greatly assisted in formulating the framework and refining the theoretical development of this work.
@@ Appendix A Synthetic Example of LoRA Deficiency
One critical limitation in the literature is the absence of convergence guarantees toward valid optimality conditions of (1). Most existing works establish convergence only with respect to the low-dimensional parameters—such as the factors $A$ and $B$ in LoRA—but do not address convergence with respect to the full-parameter matrix $W$. For instance, Jiang et al. (2024) shows that $\nabla_{A}\ell(W_{0}+A_{k}B_{k})$ and $\nabla_{B}\ell(W_{0}+A_{k}B_{k})$ vanish as $k\to\infty$, while leaving the behavior of $\nabla_{W}\ell$ uncharacterized. This gap is not merely technical: it highlights a fundamental deficiency of PEFT methods compared to standard full-parameter training. In fact, the optimal loss attained by LoRA can be arbitrarily worse than the true optimal loss of (1). To illustrate this, consider the following simple synthetic example in matrix optimization:
@@ Appendix B Review on Subspace Minimization
It is worth noting that Algorithm 1 is essentially equivalent to the classical two-loop subspace minimization scheme (Conn et al., 1994), summarized in Algorithm 4.
@@
1:Initialization $W_{0}\in\mathbb{R}^{m\times n}$, $\xi_{0}\in\mathbb{R}^{d}$, $\mathcal{M}_{0}$, an algorithmic subroutine UpdateM, an optimizer Opt, frequency $K$.
@@
4:  Update the subspace $(\mathcal{M}_{k},\widetilde{W}_{k})\leftarrow\texttt{UpdateM}(\mathcal{M}_{k-1},\widetilde{W}_{k-1})$.
@@ Appendix C Projected Subspace and Memory Efficiency
One-sided projected subspaces in PESO can offer stronger memory efficiency than LoRA. This idea is exemplified by GaLore (Zhao et al., 2024), which we now place in the PESO framework. GaLore requires memory of order $mn+mr+2nr$ (assuming $m\leq n$), compared to LoRA’s $mn+3mr+3nr$.
@@
which matches the classical projected subspace descent step (Kozak et al., 2019).
@@ Appendix D Implementation of PESO-LoRA-R
This section details implementation strategies for PESO-LoRA-R that are critical for practical stability, specifically regarding subspace smoothing and the alignment of optimizer states. While these techniques collectively enhance robustness, our empirical observations indicate that velocity alignment is the most significant factor in stabilizing training dynamics during restarts. Consequently, for a streamlined implementation that faithfully recovers our results, we strongly recommend prioritizing the velocity alignment mechanism described in Section D.2.
@@ D.1 Smoothing the Subspaces
To mitigate this, we adopt an Exponential Moving Average (EMA) of old and new subspaces, similar in spirit to how Adam (Kingma and Ba, 2014) stabilizes noisy updates. However, this is nontrivial in PESO-LoRA-R (Algorithm 2), since the pre-restart adapters $(A_{k},B_{k})$—evolved through Adam dynamics—may differ significantly in scale and coordinates from the restarted pair $(U_{k}\sqrt{\Lambda_{k}},\sqrt{\Lambda_{k}}V_{k})$ obtained from rank-$r$ SVD. A naive EMA would mismatch these terms and discard valuable exploration information.
@@
with parameter $\tau_{2}$. Since this merges the scaling of $A$ and $B$, we refactorize $S_{\mathrm{new}}$ using polar decomposition (Glentis et al., 2025):
@@ Appendix E Detailed Version of Algorithms
In this section, we provide detailed versions of PESO-LoRA-R (Algorithm 2) and PESO-LoRA-T (Algorithm 3) which are implemented in the experiments of Section 4.
Algorithm 5 PESO-LoRA-R: PESO with LoRA and Subspace ExploRation (Practical Version)
1:Pre-trained parameters $W_{0}\in\mathbb{R}^{m\times n}$, frequency $K$, scale parameter $\gamma$, learning rate $\eta$, AdamW hyperparameters $(\beta_{1},\beta_{2},\epsilon,\lambda)$.
@@
17:  $A_{k}\leftarrow A_{k-1}-\eta\lambda A_{k-1}-\eta\,\widehat{m}_{A}/(\sqrt{\widehat{v}_{A}}+\epsilon)$.
18:  $B_{k}\leftarrow B_{k-1}-\eta\lambda B_{k-1}-\eta\,\widehat{m}_{B}/(\sqrt{\widehat{v}_{B}}+\epsilon)$.
@@
Algorithm 6 PESO-LoRA-T: PESO with LoRA and Subspace ExploiTation (Practical Version)
1:Pretrained weights $W_{0}\in\mathbb{R}^{m\times n}$, initial subspace matrices $U_{0}\in\mathbb{R}^{m\times r}$, $V_{0}\in\mathbb{R}^{r\times n}$, initial coordinate $\xi_{0}\in\mathbb{R}^{r}$, frequency $K$, learning rate $\eta$, AdamW hyperparameters $(\beta_{1},\beta_{2},\epsilon,\lambda)$.
@@
13:   $U_{k}\leftarrow U_{k-1}-\eta\lambda U_{k-1}-\eta\,\widehat{m}_{U}/(\sqrt{\widehat{v}_{U}}+\epsilon)$.
14:   $V_{k}\leftarrow V_{k-1}-\eta\lambda V_{k-1}-\eta\,\widehat{m}_{V}/(\sqrt{\widehat{v}_{V}}+\epsilon)$.
@@
19:  $\xi_{k}\leftarrow\xi_{k-1}-\eta\lambda\xi_{k-1}-\eta\,\widehat{m}_{\xi}/(\sqrt{\widehat{v}_{\xi}}+\epsilon)$.
@@ F.1 Natural Language Understanding: Hyperparameter Settings
In Section 4.1, we present the results of our methods and various LoRA-based algorithms on natural language understanding tasks, following the prompt tuning configuration of (Wang et al., 2024a). The general hyperparameter settings are kept consistent across all algorithms which are shown in Table 6. To ensure a fair comparison, we follow (Zhang et al., 2025b) and tune the learning rates via grid search over $\{1\times 10^{-4},2\times 10^{-4},5\times 10^{-4},1\times 10^{-3}\}$. Additionally, following the choices of Zhang et al. (2025b), the scale parameters for LoRA-One are set to be {128, 16, 128, 128, 64} for {MNLI, SST-2, CoLA, QNLI, MRPC}.
@@ F.2.1 Comparison with LoRA-based methods
For natural language generation tasks in Section 4.2, we follow the configuration of prompt tuning and strategy of hyperparameter tuning in (Zhang et al., 2025b) to ensure fair comparison. We search the best learning rate over $\{5\times 10^{-4},\,2\times 10^{-4},\,1\times 10^{-4},\,5\times 10^{-5},\,2\times 10^{-5},\,1\times 10^{-5}\}$, and the general hyperparameter setting is summarized in Table 8. Additionally, following the choice of Zhang et al. (2025b), the scale parameters are tuned within $\{16,32,64,128\}$ for LoRA-GA, LoRA-One and our methods to achieve the best performances.
@@ F.2.2 Comparison with pretraining-oriented subspace methods
The results are shown in Table 5, demonstrating that our PEFT-specialized method PESO-LoRA-R consistently outperforms these pretraining-oriented baselines in the fine-tuning regime.
@@ F.4 Multi-Epoch Low-Rank Analysis: Hyperparameter Settings
Appendix G Asymptotic Convergence
In this section, we analyze the asymptotic convergence of the general PESO framework (Algorithm 1). To maintain generality, we do not restrict the specific choices of subspaces $\{\mathcal{S}_{k}\}$ or the subroutines Opt and UpdateM. Consequently, the derived convergence bounds may contain a bias term driven by the alignment quality between the subspace and the true gradient. We formalize this quality by assuming the subspace $\mathcal{S}_{k}$ sufficiently captures the gradient $G_{k}$ at restart indices.
Assumption 5.
There exists a sequence $\{\delta_{k}\}\geq 0$ such that $dist(G_{k},\mathcal{S}_{k})\leq\delta_{k}$ for $k$ where full gradient restart is implemented. Furthermore, $\lim_{k\to\infty}\delta_{k}<\infty$.
Remark 2.
In the general case, if $\liminf_{k\to\infty}\delta_{k}>0$, the convergence established in this section may contain a bias term depending on the approximation error. However, this bias can be eliminated in specific instantiations: (i) if $\mathcal{S}_{k}$ is chosen such that $G_{k}\in\mathcal{S}_{k}^{\text{FG}}$, then $\delta_{k}=0$ and the optimality measure converges to zero. (ii) in a deterministic setting, if $\mathcal{S}_{k}$ is chosen as the top-$r$ SVD subspace of $G_{k}$ (as in PESO-LoRA-R), the bias term vanishes. In this case, we establish exact asymptotic convergence, i.e., $\liminf_{k\to\infty}\|G_{k}\|=0$. We provide the full details in Appendix G.2. (iii) Even in the stochastic case, exact convergence can be recovered. By specifying the subspace construction as top-$r$ SVD (to control projection bias) and adopting a carefully tuned hyperparameter schedule, PESO-LoRA-R achieves exact convergence. The proof of this result (Theorem 5.1) is provided in Appendix H.
@@ G.1 Deterministic Case
For completeness, we first consider the deterministic setting, where the gradients $G_{k}$ used by PESO (Algorithm 1) are exact and free of stochastic noise. We begin by stating a mild assumption on the behavior of Opt and UpdateM, which ensures that each iteration produces a descent step.
@@
Opt and UpdateM generate the updates satisfying $\ell(W_{k})\leq\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))$ and $\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))\leq\ell(W_{k-1})$ for all $k=1,2,\cdots$.
Assumption 6 requires that both Opt and UpdateM are descent-preserving, ensuring that the objective value is non-increasing across iterations. This condition is standard and holds, for example, when both procedures are implemented via gradient descent with step size $\alpha\leq 1/L$, where $L$ is the Lipschitz constant of the gradient as specified in Assumption 1. In particular, UpdateM can be implemented as a warm-started gradient step on the subspace parameters with respect to the original loss function, initialized from the previous iterate. Such descent guarantees follow from classical smooth optimization results; see, e.g., Nesterov (2018).
@@
Let Opt and UpdateM be gradient descent schemes on $\ell$ with constant learning rate $\alpha\leq 1/L$. Then Assumption 6 holds.
@@
Suppose Assumptions 1, 5, and 6 hold. With full gradient restart enabled and learning rate $\eta_{k}=\tfrac{1}{L}$, the iterates $\{W_{k}\}$ generated by Algorithm 1 satisfy $\liminf_{k\to\infty}\|G_{k}\|\leq\lim_{k\to\infty}\delta_{k}$.
@@
Performing a full gradient restart as UpdateM, as discussed in (8), is thus equivalent to moving from $W_{k-1}$ to $\widehat{W}_{k}$. It follows that
@@
Importantly, (15) remains valid regardless of how frequently other types of UpdateM are applied between full gradient restarts, since all updates preserve the descent property by Assumption 6. In particular, whenever $\mathcal{M}_{k}$ is updated without a full gradient restart, we have